OpenStudy (anonymous):

Simplify the radical: Sqrt 50 2 sqrt 5 ?

OpenStudy (eyust707):

what times what equals 50?

OpenStudy (anonymous):

10*5

OpenStudy (eyust707):

yep what else?

OpenStudy (anonymous):

\[5\sqrt{10}\]

OpenStudy (eyust707):

nope

OpenStudy (anonymous):

needs to be simplified more ?

OpenStudy (eyust707):

what times what equals 50? (besides 5 *10)

OpenStudy (anonymous):

2*25

OpenStudy (eyust707):

yes! and check this out

OpenStudy (eyust707):

\[\sqrt{50} = \sqrt{25 * 2}\]

OpenStudy (eyust707):

\[= \sqrt{25} \sqrt{2}\]

OpenStudy (anonymous):

ok

OpenStudy (eyust707):

so what do you get?

OpenStudy (anonymous):

so i think my answer is gonna be . ..

OpenStudy (anonymous):

\[5\sqrt{2}\]

OpenStudy (eyust707):

=D